# Nowcast3D: Reliable precipitation nowcasting via gray-box learning

### v3

@@ Abstract
Extreme-precipitation nowcasting requires high spatial and temporal resolution together with extended lead times, yet current approaches remain constrained. Numerical weather prediction systems and their deep-learning emulators operate at relatively coarse space–time resolution and struggle to capture rapidly evolving convective systems. Radar extrapolation methods, which advect recent fields using estimated motion, have difficulty capturing the complex evolution of precipitation. Purely data-driven models, often produce overly smoothed reflectivity fields and underestimate intensity. Hybrid 2D radar-based methods discard crucial vertical information, preventing accurate reconstruction of height-dependent dynamics. We introduce Nowcast3D, a gray-box, fully three-dimensional nowcasting framework that operates directly on volumetric radar reflectivity and couples physically constrained neural operators with data-driven learning. The model learns three fields that govern reflectivity evolution: a three-dimensional flow field for advective transport, a spatially varying diffusion field for local dispersive spreading, and a residual source term for unresolved microphysical effects. These learned operators advance the forecast in time under explicit physical constraints, while a conditional diffusion model, conditioned on both the observations and the physics-based forecast, generates ensembles of future radar volumes that quantify forecast uncertainty. In a blind evaluation by 160 meteorologists, Nowcast3D is preferred in 57% of post-hoc and 51% of prior assessments. By explicitly embedding three-dimensional dynamics and uncertainty into a single framework, Nowcast3D offers a scalable and robust approach for reliable nowcasting of extreme precipitation.
Keywords: Deep Learning, Nowcasting, 3D volumetric radar reflectivity, Physics-guided
@@ Introduction
On 16 June 2024, extreme rainfall struck multiple locations in Meizhou, Guangdong Province. The resulting floods and landslides in Pingyuan County, Jiaoling County and Meixian District drove several rivers above historical levels, affected more than 160,000 people across eight administrative regions, and left 55 individuals dead or missing. Such episodes starkly illustrate the threat that intense precipitation poses to life and infrastructure, and they underscore the urgent need for nowcasting systems that provide both longer lead times and higher spatial and temporal detail to enable timely warnings and effective risk mitigation.
@@
Accurate short-term precipitation forecasts are essential for disaster preparedness, yet intense rainfall evolves rapidly and spans a wide range of spatial and temporal scales, which places severe demands on both traditional numerical models and emerging data-driven methods [zhang2023skilful]. Weather radar provides an observational foundation that is well suited to this challenge. Modern radar networks acquire three-dimensional reflectivity fields with horizontal resolution approaching 0.01° ($\approx$ 1 km), minute-scale temporal sampling and vertical spacing of a few hundred metres, yielding dense volumetric measurements of precipitating systems [starzec2018using, franch2020taasrad19, otsuka2016precipitation]. Radar reflectivity is closely related to instantaneous precipitation intensity. Time sequences of three-dimensional radar data therefore record both how precipitation patterns move and how reflectivity changes with height. Variations in echo-top height, bright-band signatures and layer-dependent reflectivity are closely linked to convective initiation, growth and propagation and strongly influence the evolution and organisation of heavy precipitation [steiner1995climatological, biggerstaff2000improved].
@@
Numerical weather prediction (NWP) remains indispensable from synoptic to subseasonal scales because it integrates the governing equations of motion and thermodynamics. However, operational NWP systems are constrained by update cycles, domain size and computational cost, which limits how often forecasts can be refreshed and at what spatial resolution they can be produced. These constraints reduce their utility for rapidly evolving convection [sun2014use]. Deep-learning efforts to accelerate or emulate NWP have improved speed and scalability [bi2023accurate, lam2023learning, kochkov2024neural, bodnar2025foundation, sun2025data, cui2025forecasting, allen2025end], but most are trained on global or regional NWP output and therefore inherit its typical grid spacing of tens to hundreds of kilometres and hourly output frequency, which falls short of convective nowcasting requirements. Complementary extrapolation techniques that infer motion from recent fields and then advect those fields forward in time, such as pySTEPS, can preserve large-scale transport and achieve good skill at very short lead times, yet positional and intensity errors grow rapidly as lead time increases [pulkkinen2019pysteps]. End-to-end data-driven models trained on large radar archives have substantially advanced spatiotemporal sequence modelling [shi2015convolutional, wang2017predrnn, gao2022earthformer, gao2022simvp, espeholt2022deep], but in the absence of explicit physical constraints their forecasts commonly smooth and weaken intense convective features, which limits their reliability in applications that depend on accurate prediction of extremes.
@@
Hybrid neural approaches that incorporate physical constraints have begun to narrow this gap [das2024hybrid]. By embedding evolution operators directly into the network architecture (for example, in NowcastNet [zhang2023skilful]), these methods tie the forecast to transport dynamics and often improve spatial coherence and realism. Important limitations nonetheless remain. Many existing hybrid methods operate on two-dimensional radar composites, most commonly the vertical maximum reflectivity (MAX(Z)), which discards information on the vertical profile and therefore cannot represent height-dependent shear, differential motion between layers or echo-top evolution that are central to many convective and frontal systems [franch2020taasrad19, starzec2018using].
To address these limitations, we introduce a unified three-dimensional nowcasting framework that operates directly on volumetric radar reflectivity and jointly models large-scale transport and small-scale variability. To mitigate near-surface sparsity in the three-dimensional reflectivity field, we apply a simple, non-neural imputation that fills only missing values and leaves all observed echoes unchanged (details in Supplementary Note 1.3). Instead of explicitly recovering a sparse three-dimensional velocity field, we represent the flow through a Helmholtz decomposition into stream-function and potential-function components [helmholtz1858integrale, cao2014application, xing2024HelmFluid], thereby providing a more spatially continuous description of motion. At each time step this velocity field prescribes a single transport direction and magnitude at every grid point; to capture local condensation and diffusive spreading that cannot be represented by pure advection alone, we learn a spatially varying anisotropic diffusion field. Semi-Lagrangian advection [staniforth1991semi, zhang2023skilful] combined with this anisotropic diffusion represents both fluid-like deformation and small-scale spreading, while remaining microphysical and convective perturbations are captured by a learned residual branch. Building on this physics-grounded deterministic core, we add two complementary conditional diffusion modules [ho2020denoising, song2020denoising, gong2024cascast, yu2024diffcast, price2025probabilistic]: one that focuses on the evolution of the large-scale organisation of the reflectivity field, and another that models residual, fine-scale variability. This decomposition provides a route to ensemble prediction and uncertainty quantification, aligning the system with operational needs for probabilistic guidance [li2024generative]. In an evaluation involving 160 meteorologists across China, more than half of the participants expressed a clear preference for our forecasts over existing alternatives, underscoring the potential of this framework to improve the reliability and usefulness of precipitation nowcasting.
@@
Figure 1: (a) Physical decomposition of reflectivity evolution. Reflectivity changes are partitioned into three contributions: advection by the 3D wind field, local diffusion representing small-scale spreading, and microphysical tendencies that alter reflectivity in situ. (b) Overall model schematic. Phy-Pred Network first infers latent physical fields from 3D radar; a physics-based predictor then evolves the three processes above. The physics forecast conditions a conditional diffusion model, and multiple independent noise samples produce an ensemble of probabilistic forecasts. (c) Deterministic physics backbone. The deterministic predictor inserts explicit advection, diffusion and microphysical operators into each temporal update. (d) The diffusion model captures physics-scale structure and stochastic residuals; varying noise trajectories under fixed conditioning produce calibrated probabilistic forecasts.
@@
We present Nowcast3D, a deep learning framework for forecasting volumetric radar reflectivity. The core idea is to decompose complex atmospheric evolution into physically interpretable components and to couple a deterministic, physics-based extrapolation with a probabilistic generative model in a single hybrid architecture.
@@
Our methodology decomposes the evolution of radar reflectivity into three dominant physical processes: advection, representing the coherent drift of hydrometeors by the three-dimensional flow; local turbulent diffusion, capturing the tendency of echoes to spread toward neighbouring grid points as a result of small-scale mixing; and a microphysical tendency that accounts for in situ mass and phase changes (Figure 1a) [sigrist2015stochastic]. In this formulation, the advective field specifies, at each grid point, a dominant direction of motion, whereas the diffusive term relaxes the assumption of a single trajectory and allows reflectivity to fan out locally rather than move rigidly along a unique path. This advection–diffusion–microphysics decomposition provides a compact and interpretable description of reflectivity dynamics and serves as a strong prior for the deterministic core of Nowcast3D. The overall architecture consists of two tightly coupled stages (Figure 1b): a physics-driven prediction module and a probabilistic generation module. The physics-driven module first produces a single, dynamically constrained deterministic forecast, which then serves as the reference state from which the generative module constructs an ensemble of plausible future evolutions, thereby capturing the intrinsic uncertainty of precipitation development. A detailed overview of the full Nowcast3D pipeline is provided in Extended Data Figure 1.
@@
The physics-driven module (Figure 1c) extrapolates the radar field by using a neural network to infer these governing processes from a sequence of past observations and a space - time embedding. In a single forward pass, an encoder processes the historical three-dimensional radar volumes and outputs complete space–time fields of (i) an advective drift velocity, (ii) a local diffusion field and (iii) a residual source term over the entire forecast horizon. Rather than explicitly reconstructing a sparse three-dimensional velocity field, one component of the network predicts scalar potential and vector stream-function fields, which are combined through a Helmholtz decomposition [helmholtz1858integrale, cao2014application, xing2024HelmFluid] to yield a more spatially continuous drift velocity field. In parallel, other components infer a spatially varying, anisotropic diffusivity tensor that governs the local spreading of echoes and a source field representing the net effect of microphysical processes. As summarised in Equation 7, the final deterministic forecast is obtained by numerically solving the advection–diffusion–source system defined by these three operators, using the last observed radar state as the initial condition and the network-inferred fields as the governing parameters. The detailed configuration of the space–time embeddings is given in Supplementary Note 1.5. A detailed description of the model architecture and implementation is provided in Supplementary Note 2.1, and an analysis of the individual physical terms, together with comparisons to variants that learn the velocity field directly, is given in Supplementary Note 5.1.
@@
The probabilistic refinement stage is implemented as a conditional diffusion model with a dual-branch architecture (Figure 1d). Both branches are conditioned on the same context, consisting of the past three-dimensional radar volumes and the physics-driven deterministic forecast, so that the generative model remains anchored to the physically constrained evolution. The first, structure branch learns to generate a full three-dimensional reflectivity field that captures the overall spatial and temporal organisation of echoes over the forecast horizon. The second, intensity branch learns a residual intensity field relative to the deterministic forecast, representing scale-localised amplification or weakening of precipitation intensity on top of the baseline solution. Using these two components, we can construct several types of samples: a purely generative structural realisation, an intensity-refined version of the deterministic forecast, and a sample that combines the generated structure with the learned intensity residual. As summarised in Equation 8, drawing multiple independent noise vectors and propagating them through both diffusion branches under the shared conditioning yields a multi-member ensemble of three-dimensional reflectivity forecasts, providing a quantification of uncertainty in both spatial organisation and intensity. A detailed description of the model architecture and implementation is provided in Supplementary Note 2.2, and examples of individual ensemble members, together with results from a variant without the probabilistic refinement module, are shown in Supplementary Note 5.2.
@@
We evaluate the predictive skill and potential operational utility of Nowcast3D against strong baselines that are representative of extrapolation, hybrid and purely data-driven nowcasting. The operational convective extrapolation system pySTEPS [pulkkinen2019pysteps] serves as a physics-informed reference widely adopted by national meteorological services. NowcastNet [zhang2023skilful] combines physics-informed priors with generative modeling and has demonstrated expert-validated operational value. SimVP [gao2022simvp], originally developed for generic video prediction, is included as a competitive deep learning baseline.
@@
All models are trained and evaluated on large-scale radar observations from China. The primary training dataset consists of fixed-length radar sequences from South China (Huanan area) during 2024, a region chosen for its diverse convective regimes. To assess cross-regional performance, we evaluate on an independent test set from North China (Huabei area), with no geographical overlap with the training domain. To further examine behaviour in high-resolution urban settings, we design a dedicated experiment over Maoming at 0.01° spatial resolution. In this experiment, Nowcast3D and all other learning-based methods are first trained on the South China dataset and then fine-tuned on a subset of 0.01° radar sequences from Maoming in 2024, while independent Maoming subset from 2025 are reserved for testing. Nowcast3D is always trained on native volumetric (3D) radar data, whereas all 2D baseline models are trained on column-maximum reflectivity derived from the same volumes to ensure a fair comparison. Across all experiments, models share identical input–output horizons and preprocessing pipelines. Hyperparameters for all learning-based models are tuned on a dedicated validation set, and the best-performing checkpoint is used for testing. No external data are used.
@@
Our evaluation framework employs three complementary metrics to jointly assess forecast quality. First, to quantify location-aware event detection, we use the neighbourhood Critical Success Index (CSI) [jolliffe2011forecast], which evaluates hits, misses and false alarms within prescribed radii at multiple reflectivity thresholds. Second, to diagnose the realism of predicted precipitation structures across spatial scales, we analyze the power spectral density (PSD) [sinclair2005empirical], comparing spectral slopes and power ratios of forecasts against observations from convective to mesoscale ranges. Third, to measure perceptual similarity and structural fidelity, we employ the Learned Perceptual Image Patch Similarity (LPIPS) metric [zhang2018unreasonable]. For a fair comparison with two-dimensional baselines, all three metrics are evaluated on column-maximum reflectivity fields, obtained by applying the same vertical projection to the three-dimensional Nowcast3D forecasts.
@@
Beyond standard reflectivity-based metrics, the unique ability of Nowcast3D to reconstruct the underlying atmospheric motion enables direct physical validation. Accordingly, we independently verify the predicted three-dimensional wind fields against co-located measurements from wind-profiler radars and ground-based weather stations across multiple altitudes.
@@
Figure 2: Nowcast3D accurately forecasts the evolution of a severe mesoscale convective system. a, Radar reflectivity of the convective system over North China at the forecast initialization time. The black box indicates the forecast domain, and red dots mark the locations of wind profiler radar stations. b, Quantitative forecast skill comparison using the CSI at 30 dBZ and 40 dBZ reflectivity thresholds over the 3 hour forecast horizon. c, Analysis of spatial scale fidelity using PSD. Forecast spectra are compared to the observed spectrum (black dashed line). d, Perceptual similarity assessment using the LPIPS metric. Lower values indicate better performance. e, Vertical wind profile validation at the Haidian station. The predicted zonal (east–west) and meridional (north–south) wind components are compared against wind profiler observations at 60 minute. f, Comparison of the predicted horizontal wind field with wind profiler observations in Beijing at an altitude of 2,500 m, 60 minutes into the forecast. g, Visual comparison of forecast reflectivity fields from Nowcast3D and baseline models against the observed radar sequence.
@@
We evaluate model performance on the extreme rainfall event that affected Beijing on 28 July 2025. This event triggered a red heavy-rain warning from the Beijing Meteorological Observatory, the highest tier in China’s warning system, and caused substantial societal impacts, including 30 reported fatalities. Our analysis focuses on the period of intense convection between 15:00 and 18:00 UTC, during which the storm evolved rapidly and developed complex internal structure (Figure 2a).
@@
The storm system consisted of two main convective cells. A western cell moved eastward and produced broad, mainly moderate rainfall. The eastern cell, situated closer to Beijing, remained the primary hazardous core with persistently high reflectivity. As the event evolved, the two systems became more organised and gradually merged, and the region of heaviest rainfall developed into a pronounced V-shaped echo. One arm of this V extended eastward, while the other stretched toward the southwest. Their motions differed: the eastern arm continued to drift eastward, whereas the southwestern arm shifted slightly southwestward and then became nearly stationary. This contrasting motion produced strong differential movement within the system and continually reshaped the echo, posing two key challenges for forecasting: capturing the evolving morphology and maintaining the extreme rainfall intensity without artificial smoothing or rapid decay.
@@
A comparison between the predicted and observed radar sequences highlights clear differences among the methods (Figure 2g). The SimVP forecast exhibits rapid loss of intensity and substantial distortion of the storm structure. pySTEPS and NowcastNet better preserve reflectivity amplitudes and sharp gradients, but show large position errors and fail to reproduce the detailed evolution of the hazardous eastern cell. In contrast, Nowcast3D closely follows the complex evolution of the system, accurately tracking the displacement of the eastern cell while retaining its internal structure and high reflectivity. These qualitative impressions are supported by quantitative metrics. Nowcast3D achieves consistently higher neighbourhood CSI scores across all lead times, indicating more accurate prediction of heavy-rainfall locations (Figure 2b). In addition, Nowcast3D outperforms all baselines in both PSD and LPIPS metrics, demonstrating more realistic small-scale variability and higher perceptual similarity to the observed radar fields (Figure 2c,d).
@@
An important feature of our framework is its ability to infer the internal three-dimensional wind field. To evaluate this physical component, we compare the predicted winds with independent observations from wind-profiler radars. Figure 2f shows the forecast horizontal wind at 2,500 m altitude over Beijing, 60 minutes into the prediction, which agrees well with co-located profiler measurements. Figure 2e further presents vertical profiles at the Haidian station. The model reproduces the overall structure of both the zonal (east–west) and meridional (north–south) wind components, with some deviation in the zonal trend but similar magnitudes, and closer agreement for the meridional component. Layer-by-layer comparisons of the three-dimensional radar forecasts are provided in Extended Data Figure 3.

### v4

@@ Abstract
Reliable nowcasting of extreme precipitation remains a major challenge, owing to the strongly nonlinear, multiscale, and nonstationary nature of convective systems in 3D space. Radar observations offer the backbone of nowcasting, yet existing methods struggle to faithfully predict extreme precipitation evolution: physics-based extrapolation falls short in capturing growth and decay, purely deterministic data-driven methods often produce over-smoothing prediction and underestimate extremes, and purely generative models usually lack physical consistency. While hybrid schemes that blend machine learning with physical principles are promising, they remain confined to 2D composite radar reflectivity. This collapses the atmosphere into a single layer and discards the critical vertical structure, thereby hindering accurate modeling of height-dependent dynamics. We introduce Nowcast3D, a gray-box, fully 3D nowcasting framework that operates directly on volumetric radar reflectivity. The end-to-end learning model integrates physically constrained neural operators (encoding advection, local diffusion, and microphysical processes) with a conditional diffusion model to generate ensemble forecasts with quantified uncertainty. Trained on provincial-scale 3D radar volumes spanning a $10.24^{\circ}\times 10.24^{\circ}$ region and fine-tuned on data in a $2.56^{\circ}\times 2.56^{\circ}$ city-scale region ($0.01^{\circ}\approx 1$ km), Nowcast3D delivers near-real-time radar-reflectivity forecasts with lead times up to 3 hours. It performs robustly and remains superior compared with several competitive baselines in both cross-region and temporal-out-of-sample evaluations. Notably, Nowcast3D is also capable of inferring wind fields without any labeled training data, supporting the physical plausibility of the learned transport dynamics. In a nationwide blind evaluation by 160 meteorologists, Nowcast3D ranked first for its competitive performance and was preferred in 57% of post-hoc assessments, surpassing the leading baseline (27%). These results highlight its potential of reliability and operational value for extreme precipitation nowcasting.
Keywords: Precipitation nowcasting, 3D volumetric radar reflectivity, Physics-guided learning
@@ Introduction
On 16 June 2024, extreme rainfall struck multiple locations in Meizhou, Guangdong Province, China. The resulting floods and landslides in Pingyuan County, Jiaoling County and Meixian District caused several rivers to exceed historically recorded levels, impacted over 160,000 people across eight administrative regions, and resulted in 55 individuals dead or missing. Such events starkly illustrate the severe threat posed by intense precipitation to lives and infrastructure. They also highlight the pressing need for nowcasting systems capable of delivering longer lead times and higher accuracy, in order to facilitate timely warnings and improve risk mitigation.
@@
Accurate short-term precipitation forecasting is critical for pre-disaster planning. However, the rapid evolution of intense rainfalls across a wide range of spatial and temporal scales poses severe challenges to both traditional numerical models and emerging data-driven methods [zhang2023skilful]. Weather radar provides an observational foundation well-suited to this task. Modern radar networks acquire three-dimensional (3D) reflectivity fields with a horizontal resolution approaching 0.01° ($\approx$ 1 km), minute-scale temporal sampling, and vertical spacing of a few hundred meters, yielding dense volumetric measurements of precipitating systems [starzec2018using, franch2020taasrad19, otsuka2016precipitation]. Radar reflectivity is closely related to instantaneous precipitation intensity. Consequently, time sequences of 3D radar data capture not only the movement of precipitation patterns but also the vertical structure of reflectivity. Variations in echo-top height, bright-band signatures, and layer-dependent reflectivity are closely linked to the initiation, growth, and propagation of convection, which strongly governs the evolution and organization of heavy precipitation [steiner1995climatological, biggerstaff2000improved].
@@
Numerical weather prediction (NWP) remains indispensable from synoptic to subseasonal scales, as it explicitly integrates the governing equations of motion and thermodynamics. However, operational NWP systems are constrained by update cycles, domain size, and computational cost, which limit the refresh rate and spatial resolution of forecasts. These limitations reduce their effectiveness for capturing rapidly evolving convection events [sun2014use]. While deep learning efforts to accelerate or emulate NWP have achieved gains in speed and scalability [bi2023accurate, lam2023learning, kochkov2024neural, bodnar2025foundation, sun2025data, cui2025forecasting, allen2025end], these models are trained on global or regional NWP data. Consequently, they typically inherit a coarse grid spacing of tens to hundreds of kilometers and hourly output frequency, which falls short of the resolution required for convective nowcasting requirements.
Radar-observation-based forecasting methods remain the dominant paradigm for precipitation nowcasting. physics-based extrapolation techniques, such as pySTEPS [pulkkinen2019pysteps], infer motion from recent fields and then advect them forward in time. Although these methods can preserve large-scale transport and achieve good skill at very short lead times, their positional and intensity errors grow rapidly as the lead time increases. End-to-end data-driven models trained on extensive radar archives have shown a great potential, e.g., DGMR [ravuri2021skilful], SimVP [gao2022simvp], among others [shi2015convolutional, wang2017predrnn, gao2022earthformer, espeholt2022deep]. However, in the absence of explicit physical constraints, their forecasts tend to over-smooth and weaken intense convective features, limiting their reliability for applications that depend on accurate prediction of extreme events.
@@
Hybrid learning approaches that embed physical processes into neural networks have emerged to narrow this gap [das2024hybrid]. For instance, NowcastNet [zhang2023skilful] unifies physical-evolution schemes and conditional-learning methods, thereby anchoring forecasts to transport dynamics and enhancing the physical coherence of prediction. However, notable limitations remain. Many existing hybrid methods operate on 2D radar composites, most commonly the vertical maximum reflectivity (e.g., MAX(Z)). This discards critical information on the vertical structure of atmosphere and therefore fails to capture features that are central to many convective and frontal systems [franch2020taasrad19, starzec2018using], such as height-dependent air shear and differential motion between layers or echo-top evolution.
To address these limitations, we develop Nowcast3D, an end-to-end gray-box nowcasting network that unifies physics-inspired neural operators and a conditional diffusion model, operating directly on 3D volumetric radar reflectivity. By modeling convective systems in 3D space, Nowcast3D explicitly resolves height-dependent the atmospheric structure and its temporal evolution. Specifically, the model learns to decompose the observed reflectivity dynamics into three physically interpretable fields: a 3D flow field that governs advective transport, a spatially varying anisotropic diffusion field that captures local dispersive spreading, and a residual source term that accounts for unresolved microphysical and convective contributions. These components form a physics-grounded yet learnable deterministic core that advances forecasts under explicit constraints. To correct for prediction errors and represent uncertainty, this core is coupled with a diffusion model, conditioned jointly on both the input radar observations and the physics-informed forecast, to produce ensemble predictions for robust uncertainty quantification.
[…]
@@
Figure 1: Schematic of the Nowcast3D model. a, Physical decomposition of reflectivity evolution. The changes in 3D radar reflectivity are decomposed into three contributions: advection by the 3D wind field, local diffusion capturing small-scale spreading, and microphysical tendencies that modify reflectivity in situ. b, Overall model architecture. The Phy-Pred Network infers latent physical fields from input 3D radar with spatiotemporal embedding, and a physics-based predictor advances the system according to the three processes in a. The resulting physics forecast conditions a diffusion generator which produces an ensemble of probabilistic forecasts via drawing multiple independent noise realizations. Details of Nowcast3D are provided in Extended Data Figure 1. c, Deterministic physics backbone (the Phy-Pred Process). The deterministic predictor performs each temporal update by explicitly applying advection, diffusion, and microphysical operators within the Phy-Pred Process module. d, Diffusion-based probabilistic generation (the Diff-Gen Network). The diffusion model learns physics-consistent mesoscale structure by the Diffusion Structure Network and stochastic residual components by the Diffusion Intensity Network. Under fixed conditioning, different sampled noise trajectories produce calibrated probabilistic 3D reflectivity forecasts.
@@
Nowcast3D is designed to forecast precipitation, in particular extreme events, based on volumetric radar reflectivity. The core idea is to decompose complex atmospheric evolution into physically interpretable components and establish a unified learning architecture that couples a physics-inspired deterministic extrapolation network with a diffusion-based generative model for probabilistic forecasts (see Figure 1).
@@
Our methodology decomposes the evolution of radar reflectivity into three dominant physical processes: advection, which captures the coherent drift of hydrometeors by 3D flow; local turbulent diffusion, representing the spreading of echoes toward neighbouring grid points due to small-scale mixing; and a microphysical tendency that accounts for in-situ mass and phase changes (Figure 1a) [sigrist2015stochastic]. In this formulation, the advective field specifies, at each grid point, a dominant direction of motion, whereas the diffusive term relaxes the assumption of a single trajectory and allows reflectivity to fan out locally rather than move rigidly along a unique path. This advection–diffusion–microphysics decomposition provides a compact and interpretable description of reflectivity dynamics and serves as a strong prior for the deterministic core of Nowcast3D. The overall architecture consists of two tightly coupled stages (Figure 1b): a physics-driven prediction module and a probabilistic generation module. The physics-driven module first produces a single, dynamically constrained deterministic forecast, which then serves as a reference state. The generative module constructs an ensemble of plausible future evolutions based on the reference, thereby capturing the intrinsic uncertainty of precipitation development. An overview of the full Nowcast3D pipeline is depicted in Extended Data Figure 1, with detailed formulations presented in the Methods section.
@@
The physics-driven module (Figure 1c) extrapolates the radar field using a neural network to infer the governing processes from a sequence of past observations and spatiotemporal embedding. In a single forward pass, an encoder processes the historical 3D radar volumes and outputs complete spatiotemporal fields of (i) an advective drift velocity, (ii) a local diffusion field, and (iii) a residual source term over the entire forecast horizon. Rather than explicitly reconstructing a sparse 3D velocity field, one branch of the network predicts scalar potential and vector stream-function fields, which are combined through a Helmholtz decomposition [helmholtz1858integrale, cao2014application, xing2024HelmFluid] to produce a more spatially continuous drift velocity field. In parallel, other network components infer a spatially varying, anisotropic diffusivity tensor that governs the local spreading of echoes, and a source field representing the net effect of microphysical processes. As summarized in Eq. (7), the final deterministic forecast is obtained by numerically solving the advection–diffusion–source system defined by these three operators, using the last observed radar state as the initial condition and the network-inferred fields as the governing parameters. The configuration of the spatiotemporal embeddings is given in Supplementary Note 1.5. A detailed description of the model architecture and implementation is provided in Supplementary Note 2.1. The analysis of the individual physical terms, together with comparisons to variants that learn the velocity field directly, is given in Supplementary Note 5.1.
@@
The probabilistic refinement stage is implemented via a conditional diffusion model with a dual-branch architecture (Figure 1d) [ho2020denoising, song2020denoising, gong2024cascast, yu2024diffcast, price2025probabilistic]. Both diffusion branches are conditioned on the same context, consisting of the past 3D radar volumes and the physics-driven deterministic forecast, so that the generative model remains anchored to the physically constrained evolution. The first branch (aka, the structure network) learns to generate a full 3D reflectivity field that captures the overall spatial and temporal organization of echoes over the forecast horizon. The second branch (aka, the intensity network) learns a residual intensity field relative to the deterministic forecast, representing scale-localized amplification or weakening of precipitation intensity on top of the baseline solution. Using these two components, we can construct several types of samples: a purely generative structural realization, an intensity-refined deterministic forecast, and a sample that combines the generated structure with the learned intensity residual. As summarized in Eq. (8), drawing multiple independent noise vectors and propagating them through both diffusion branches under the shared conditioning yields a multi-member ensemble of 3D reflectivity forecasts. This provides a quantification of uncertainties in both spatial organization and intensity [li2024generative]. A detailed description of the model architecture and implementation is provided in Supplementary Note 2.2. Examples of individual ensemble members, together with results from an ablation variant without the probabilistic refinement module, are shown in Supplementary Note 5.2.
@@
We evaluate the predictive skill and potential operational utility of Nowcast3D against three types of competitive baselines including extrapolation-based, purely data-driven, and hybrid nowcasting models (pySTEPS [pulkkinen2019pysteps], SimVP [gao2022simvp], and NowcastNet [zhang2023skilful], respectively). Specifically, the operational convective extrapolation approach pySTEPS [pulkkinen2019pysteps] serves as a physics-informed reference widely used in national meteorological services. SimVP [gao2022simvp], a deep learning model originally developed for generic video prediction, is considered as a competitive purely data-driven baseline. NowcastNet [zhang2023skilful] is included as a hybrid baseline, which integrates physics-informed priors with generative modeling and has demonstrated expert-validated operational value.
@@
All models are trained and evaluated on large-scale radar observations in China. The primary training dataset consists of fixed-length radar sequences covering a $10.24^{\circ}\times 10.24^{\circ}$ region at $0.04^{\circ}$ resolution in South China (Huanan area) during 2024, a region chosen for its diverse convective regimes. To assess cross-regional generalization, we evaluate the models on an independent test dataset in North China (Huabei area), having no geographical overlap with the training domain. To further examine the performance in high-resolution urban settings, we design a dedicated experiment in a city-scale $2.56^{\circ}\times 2.56^{\circ}$ region (Maoming, Guangdong Province, China) at $0.01^{\circ}$ resolution. In this experiment, Nowcast3D and all other learning-based methods are first pretrained on the South China dataset and then fine-tuned on a subset of 0.01° radar sequences collected in Maoming during 2024, while an independent 2025 Maoming subset held out for testing. Note that Nowcast3D is directly trained on native volumetric (3D) radar data, whereas all the baseline models are trained on 2D column-maximum reflectivity derived from the same volumes to ensure a fair comparison. Across all the experiments, models share identical input–output horizons and preprocessing pipelines. Hyperparameters for all learning-based models are tuned on a dedicated validation set, and the best-performing checkpoint is used for final testing. No other external data are used.
@@
Our evaluation framework employs three complementary metrics to jointly assess the forecast quality. First, to quantify location-aware event detection, we use the neighborhood Critical Success Index (CSI) [jolliffe2011forecast], which evaluates hits, misses and false alarms within prescribed radii at multiple reflectivity thresholds. Second, to evaluate the realism of predicted precipitation structures across spatial scales, we analyze the power spectral density (PSD) [sinclair2005empirical], comparing spectral slopes and power ratios of forecasts against observations from convective to mesoscale ranges. Third, to measure perceptual similarity and structural fidelity, we employ the Learned Perceptual Image Patch Similarity (LPIPS) metric [zhang2018unreasonable]. For a fair comparison with the 2D baselines, all these three metrics are evaluated on column-maximum reflectivity fields, obtained by applying the same vertical projection to the three-dimensional Nowcast3D forecasts.
@@
It is noted that, beyond standard reflectivity-based metrics, the unique capability of Nowcast3D to reconstruct the underlying atmospheric motion enables direct physical validation. We therefore verify its predicted 3D wind fields against co-located measurements by wind-profiler radars and ground-based weather stations at multiple altitudes.
@@
Figure 2: Nowcast3D accurately forecasts the evolution of a severe mesoscale convective system ($T$ = 15:00:$0$ pm, July 28, 2025). a, Radar reflectivity of the convective system over North China at the forecast initialization time. The black box indicates the forecast domain, and red dots mark the locations of wind-profiler radar stations. b, Quantitative comparison of forecast skills in context of CSI at 30 dBZ and 40 dBZ reflectivity thresholds over a 3-hour forecast horizon. c, Analysis of spatial scale fidelity based on PSD. The forecast spectra are compared to the observed spectrum (black dashed line). d, Perceptual similarity assessment based on the LPIPS metric. Lower values indicate better performance. e, Vertical wind profile validation at a station located in Haidian district, Beijing. The predicted zonal (east–west) and meridional (north–south) wind components are compared with wind profiler observations at the 60-min lead time. f, Comparison of the predicted horizontal wind field with wind-profiler observations in Beijing at 2,500 m altitude (60-min lead time forecast). g, Forecast reflectivity fields by Nowcast3D and baseline models, compared with the observed radar sequence. The frame-by-frame comparison at 6-min intervals is provided in Supplementary Video 1.
@@
We evaluate our model performance on the extreme precipitation event that affected Beijing, China on July 28, 2025. This event triggered a red heavy-rain warning (the highest level in China’s warning system) from the Beijing Meteorological Observatory, and resulted in substantial societal impacts, including 30 reported fatalities. Our analysis focuses on the period of most intense convection, from 15:00 to 18:00 UTC, during which the storm evolved rapidly and developed complex internal structures (Figure 2a).
@@
The storm system comprised two main convective cells. A western cell moved eastward, producing broad, mainly moderate rainfalls. The eastern cell, situated closer to Beijing, remained the primary hazardous core with persistently high reflectivity. Over time, the two systems became more organized and gradually merged, with the heaviest rainfall region evolving into a pronounced V-shaped echo. One arm of this V extended eastward, while the other stretched southwest. Their motions differed: the eastern arm continued drifting eastward, whereas the southwestern arm shifted slightly southwest before becoming nearly stationary. This contrasting motion created strong differential movement within the system, continually reshaping the echo and posing two key challenges for forecasting: capturing the evolving morphology, and maintaining extreme rainfall intensity without artificial smoothing or rapid decay.
@@
A comparison of the predicted and observed radar sequences highlights clear differences among different methods (Figure 2g). The SimVP forecast exhibits rapid loss of intensity and considerable distortion of the storm structure. pySTEPS and NowcastNet better preserve reflectivity amplitudes and sharp gradients, but exhibit large position errors and fail to reproduce the detailed evolution of the hazardous eastern cell. In contrast, Nowcast3D closely follows the complex evolution of the system, accurately tracking the displacement of the eastern cell while maintaining its internal structure and high reflectivity. These qualitative findings are supported by quantitative metrics. Nowcast3D achieves consistently higher neighborhood CSI scores across all lead times, indicating more accurate prediction of heavy-rainfall locations (Figure 2b). Nowcast3D also outperforms all the baselines in both PSD and LPIPS metrics, demonstrating more realistic small-scale variability and higher perceptual similarity to the observed radar fields (Figure 2c,d).
@@
An important feature of our framework is its ability to infer the internal 3D wind field. To evaluate this physical component, we compare the predicted winds with independent observations from wind-profiler radars. Figure 2f shows the predicted horizontal wind field at 2,500 m altitude over Beijing a 60‑min lead time, which aligns with co-located profiler measurements. Figure 2e further displays the vertical wind profiles at a station located in Haidian district, Beijing. The model reproduces the overall structure of both the zonal (east–west) and meridional (north–south) wind components capturing similar magnitudes throughout. While some deviation is seen in the zonal trend but similar magnitudes, the meridional component shows closer agreement. Layer-by-layer comparisons of the 3D radar reflectivity forecasts are provided in Extended Data Figure 2.